### v1

# WorldPrompter: Traversable Text-to-Scene Generation

Abstract
Scene-level 3D generation is a challenging research topic, with most existing methods generating only partial scenes and offering limited navigational freedom. We introduce WorldPrompter, a novel generative pipeline for synthesizing traversable 3D scenes from text prompts. We leverage panoramic videos as an intermediate representation to model the 360° details of a scene. WorldPrompter incorporates a conditional 360° panoramic video generator, capable of producing a 128-frame video that simulates a person walking through and capturing a virtual environment. The resulting video is then reconstructed as Gaussian splats by a fast feedforward 3D reconstructor, enabling a true walkable experience within the 3D scene. Experiments demonstrate that our panoramic video generation model achieves convincing view consistency across frames, enabling high-quality panoramic Gaussian splat reconstruction and facilitating traversal over an area of the scene. Qualitative and quantitative results also show it outperforms the state-of-the-art 360° video generators and 3D scene generation models.
Figure 1: We present WorldPrompter, a method that generates traversable environments from a text prompt alone (top). Our method first generates a spatially coherent video (2nd row) that travels through the environment described by the prompt. Next, we elevate the scene into a 3D Gaussian Splat (bottom), from which we project some views into standard perspective images (3rd row).
*Work done as an intern at Adobe Research.
1 Introduction
Scene-level 3D generation has attracted increasing research interest in recent years. For example, methods like RealmDreamer [38] and LucidDreamer [10] can generate Gaussian Splatting (GS) 3D scenes from text prompts. They rely on single-view diffusion models and monocular depth estimation to generate scenes progressively from one view to another by 2D inpainting. However, the 3D scenes created by these methods are confined to a limited viewable area and are not freely traversable. This constraint differentiates these approaches from a ”complete” scene generation that offers a navigable virtual world, similar to what is found in interactive video games.
In this paper, we propose WorldPrompter, the first generative pipeline that produces a navigable 3D scene from text prompts. Training a model to directly synthesize a full scene can be very challenging due to highly scarce scene-level 3D data. Drawing inspiration from conventional 3D capture methods and recent advancements in video generation models ([32, 54, 4]), we propose using videos as an intermediate representation to model complete 3D environments. Specifically, we train a diffusion model to generate videos traversing the environment based on text prompts. Moreover, instead of synthesizing captures from perspective cameras, we leverage panoramic cameras that directly capture the entire surroundings of an environment, providing 360° viewpoints that are ideal for our application.
Specifically, WorldPrompter is a two-stage pipeline composed of a 360° panoramic video generation model and a 3D reconstruction step to generate the final 3D Gaussians, a summary of which is shown in Fig. 1. The text-conditioned panoramic video generation model is trained to generate a 360° panoramic video that simulates a person walking through an environment while holding a panoramic camera to capture the scene. To train the video generator, we curated a dataset comprising approximately 1,700 panoramic videos and a panorama image dataset containing more than 200,000 high-quality panoramic in-the-wild captures. To enhance the generative capability of our method, we fine-tuned a Diffusion Transformer (DiT)-based video generator on our dataset to produce 10.6-second 360° panoramic videos (128 frames at 12 fps). We introduce a masked diffusion loss to ensure a clean panoramic capture sequence free from the person holding the camera, ready for subsequent reconstruction purposes. From the generated videos, we extract panoramic frames and project them from their original spherical coordinates onto the image plane to obtain perspective images from multiple viewpoints throughout the scene. The reconstruction step then builds the final 3D Gaussian Splat (3DGS) scene from these cropped images using a fast feedforward reconstruction model [56]. The resulting scene allows users to navigate freely within a virtual environment.
We show that our 360° video generation model achieves state-of-the-art generation capability in view coverage, temporal consistency, and visual quality. The high view consistency enables the videos to be reliably used as proxies for 3D scene generation, resulting in a significantly better view coverage of the generated scene and a traversable scene, more flexible than prior art. We highlight our contributions as follows:
  • We present WorldPrompter, the first generative pipeline to synthesize a complete 3D scene from a text prompt.
  • We propose leveraging panoramic capture videos as intermediate proxies for full 3D scene generation, addressing the scarcity of complete 3D scene datasets.
  • We introduce a high-quality 360° panoramic video generation model trained with mixed datasets and a masked diffusion loss achieving superior temporal consistency and view coverage compared to previous work.
  • We show that our 360° video is consistent enough to be efficiently turned into a 3DGS scene.
Figure 2: Overview of WorldPrompter. (top) We train a text-to-video model on a mix of 360° videos and images depicting in-the-wild environments. As it is challenging to avoid the person and camera equipment being visible in the video capture, we mask these elements out of the frame using a pretrained image segmentation model [1], and obtain the prompts from the video frames using LLaVA [24] for the videos, and BLIP-2 [19] for images. (bottom) At inference time, a user supplies a text prompt to our text-to-video model, which produces a “walk-through” video of the scene, which we reconstruct into a 3D Gaussian splat representation using Long-LRM [57].
2 Related Work
2.1 Panoramic Image and Video Generation
In computer graphics, 360° panoramas—often called IBLs or environment maps—are widely used to represent distant lighting and environments [12]. Numerous studies estimate lighting as High Dynamic Range (HDR) panoramas from input photos [51, 44, 11]. However, the resolution of these panoramas is often limited, making them suitable for light representation but insufficient for 3D reconstruction.
Advancements in diffusion models have made high-resolution image synthesis possible [35, 13]. Panorama generation also benefits from the use of diffusion models. For example, Text2Light [9] generates 4K HDR panoramas via text-driven low-resolution generation and super-resolution inverse tone mapping. MVDiffusion [39] creates panoramas by stitching consistent multi-view images generated from text prompts using attention layers. PanFusion [52] employs a dual-branch diffusion model with equirectangular-Perspective Projection Attention (EPPA) to improve consistency and control across panorama and perspective domains.
Diffusion models have also demonstrated their potential in video generation [6, 5]. Diffusion Transformers [31] offer a more versatile method for modeling high-dimensional data, making them particularly well-suited for video generation [7]. For 360° video generation, Wang et al. propose 360DVD for generating 360° panoramic videos from text prompts and motion conditions, where they leverage a lightweight 360-Adapter to finetune the Stable Diffusion [35] text-to-image model. However, their model is limited to generating only 16 frames, whereas our 360° video generation model produces significantly more frames with superior temporal consistency.
2.2 3D Neural Representation and Reconstruction
3D reconstruction and representation is an extensively studied field. While mesh-based methods [17] and point-based methods [33] have been extensively explored, they are limited in the quality of their results. In contrast, neural representations offer a flexible way to reconstruct 3D data. For instance, Neural Radiance Fields (NeRF) [26] introduces a novel way to represent 3D scenes by encoding volumetric scene information as a continuous function parameterized by a neural network, which inspired many follow-up works [27, 2, 3] that optimize efficiency, quality, and rendering speed. However, the MLP-based paradigm requires training a distinct model per scene, making reconstruction time-consuming. Additionally, rendering views involves model queries, creating a bottleneck for real-time applications.
3D Gaussian Splatting (3DGS) [18] introduces a memory-efficient and low-overhead 3D implicit representation using Gaussians derived from Structure-from-Motion (SfM) points. By combining anisotropic covariance optimization with a fast visibility-aware rendering algorithm, 3DGS enables efficient scene representation, high-quality real-time rendering, and has quickly emerged as a new paradigm in neural reconstruction [49, 47, 16, 14, 8]. Nevertheless, 3DGS still requires an optimization process, which can be costly for larger scenes. Recent advances such as Large Reconstruction Models (LRMs) [56, 45, 48, 53, 41] target scene reconstruction with a single feed-forward evaluation during inference. They achieve a reconstruction quality comparable to optimization-based methods while significantly improving their efficiency. In our scene generation pipeline, we also utilize this state-of-the-art feed-forward reconstruction approach for long-sequence reconstruction, achieving fast reconstruction.
2.3 3D Scene Generation
3D generation has emerged as a prominent focus in generative modeling. Given the scarcity of high-quality 3D data, recent works usually leverage intermediate representations, such as multi-view images, to improve the quality and efficiency of their reconstructions. Many studies have explored generating 3D objects in various formats, including Gaussian splats and meshes. For instance, Instant3D [20] accelerates 3D asset creation from text prompts by combining sparse-view generation with transformer-based reconstruction. Similarly, MVDream [37] improves consistency through multi-view diffusion by integrating 2D and 3D data. One-2-3-45 [25] simplifies single-image 3D reconstruction, enabling view-consistent mesh generation.
Scene-level generation, however, is a more challenging task because it requires more observations to achieve comprehensive view coverage. Typically, 3D scene generation methods employ progressive pipelines [29, 10, 15]: anchor cameras are first selected to generate text-aligned 2D images and corresponding depth maps, followed by a neural reconstruction process to build the initial 3D structure. This process is iteratively repeated to expand scene coverage by generating additional camera views and inpainting missing regions. However, such methods often suffer from prolonged generation times and incomplete coverage. Other studies have explored interactive and layout-guided approaches for 3D scene generation. Systems like WonderWorld [50] provide user-friendly interfaces for scene customization, while Director3D [22] facilitates real-world scene creation with adaptive camera trajectories.
The closest method to ours is DreamScene360 [55], which generates 3D scenes from text prompts using a single panorama image, along with depth estimation and Gaussian Splats optimization. However, while the resulting scene supports 360° viewing, it suffers from a limited movable range. 4K4DGen [21], introduces panorama-to-4D generation but still remains limited in terms of traversable area. More recently, ODIN [40] proposed learning 3D from 360° videos. However, their approach extracts only cropped views from videos to perform image-to-image style novel view synthesis, under-utilizing the full potential of 360° videos. As a result, ODIN fails to model a complete 3D world and can only generate partial scenes. In contrast to these approaches, our model generates complete scenes by fully leveraging the 360° representation, ensuring consistency and comprehensive coverage through a panoramic video generation model.
3 WorldPrompter
3.1 Overview
Our WorldPrompter synthesizes a traversable, 360° viewable 3D scene in two stages: generation and reconstruction. The generation stage leverages a text-conditioned 360° panoramic video diffusion model to generate a 10.6-second panoramic capture video for a target scene, simulating someone holding the camera and walking through the environment. Later, the reconstruction stage runs COLMAP [36] to estimate precise camera poses from perspective crops of the generated videos and builds a navigable 3DGS using Long-LRM [56] with the calibrated camera poses. The generated panoramic video features rich details and geometrically consistent visual details from multiple viewpoints along the traversal trajectory, ensuring high-quality reconstruction by the reconstruction step and yielding appealing novel views in the traversable scene.
3.2 Video Generator for 360° Panoramic Captures
Training a generative model for 360° videos is challenging due to limited data; therefore, we start with an existing video generator. We leverage a pretrained text-to-video diffusion model (similar to [7]) as backbone, and fine-tune it using panoramic data.
Specifically, our model’s backbone architecture $f_{\theta}$ is based on Diffusion Transformers [31]. Typically, in DiT models, a video sequence $x\in\mathbb{R}^{T\times 3\times H\times W}$ is first tokenized into patches, yielding a set of visual tokens $\hat{x}\in\mathbb{R}^{K\times D}$, where $T$, $H$, and $W$ correspond to the temporal and spatial dimensions of the video, $K$ represents the total number of tokens, and $D$ denotes the feature dimension. Being a text-conditioned model, it requires text prompts $c$ that we encode using a pre-trained model [34] and map into the same feature space $D$ by an embedding layer. The transformer $f_{\theta}(\hat{x}_{t};c,t)$ operates on these inputs, denoising the tokens at each timestep $t$. The denoised tokens $\hat{x}_{0}\in\mathbb{R}^{K\times D}$ are subsequently decoded and reassembled into the video data $x_{0}\in\mathbb{R}^{T\times 3\times H\times W}$ via linear layers. To reduce the computational complexity caused by the high-dimensionality of video data, the model performs its denoising in latent space after encoding by a Variational AutoEncoder (VAE) [31, 35].
The scarcity of 360° video data presents a significant challenge in training this model. To develop our high-quality 360° video generator, we curate two carefully selected datasets: (1) a 360° video dataset capturing static indoor environments and (2) a 360° image dataset featuring diverse real-world scenes.
3.2.1 360° video dataset
We collect a 360° video dataset with approximately 1,700 panoramic videos. Each video is captured by a person holding a Ricoh Theta Z1 camera walking inside a nearly static indoor scene to ensure the quality of 3D scene reconstruction. Each video has around 50 seconds of capture, yielding roughly 1,500 frames. To build a dataset for training, we segment the longer video sequences into smaller chunks, typically a 128-frame video clip (5.3/10.6 second at 24/12fps). We use LLaVA [23] to generate detailed captions describing each scene.
3.2.2 360° image dataset
The 360° video data is primarily captured indoors to ensure a static environment with no moving objects during the capture process. However, this may bias the fine-tuned model, limiting its ability to generalize effectively to outdoor environments. To mitigate this issue, we gather another dataset with around 200,000 panoramic images taken with cameras on a tripod, without the photographer in the scene. We caption each image with Blip-2 [19] to describe the whole scene in detail. These stationary images, captured in-the-wild, represent a much broader variety of scenes, thereby enhancing the generalization capability of our video diffusion model.
3.2.3 Mixed Training with Masked Loss
We combine the two described datasets to train our model, enabling the text-to-360° video generator to produce both 360° videos and images, with images treated as single-frame videos. While panoramic images depict a clean environment without the photographer present, the 360° video capture often includes the photographer visible in the scene. This causes the trained model to produce camera operators in its outputs (see Fig. 7), complicating subsequent reconstruction steps. Hence, when training our diffusion models on video data, we introduce a masked diffusion loss, where the photographer is excluded from the loss computation.
To achieve this, we preprocess all 360° videos by masking out the photographer using a pretrained image segmentation model [1] on each frame in a video. For a 128-frame chunk, we merge all segmentation masks to minimize segmentation errors and ensure a clean region without the person. Additionally, we conservatively mask out the bottom region of the equirectangular panoramic frames to eliminate visible camera equipment and the photographer’s hand. Assuming the corresponding binary mask $M$ is resized into the latent space by nearest neighbor interpolation, the masked diffusion loss is computed as
$\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1)}\left[\left\lVert M\odot\left(\epsilon_{t}-f_{\theta}(\hat{x}_{t};c,t)\right)\right\rVert^{2}\right],$ (1)
where $\epsilon_{t}$ is the sampled noise at timestep $t$, and $M$ has the same dimensions as $\epsilon_{t}$, where $M[i,j]=1$ includes the region in the loss calculation, and $M[i,j]=0$ excludes it (indicating the visible person).
Our video generator is finally trained on a mix of $1/3$ image data and $2/3$ video data, achieving an empirically reasonable balance. The video diffusion model generates a 128-frame video clip at 12 fps, equivalent to a 10.6-second 360° capture video at a native resolution of 352$\times$704 pixels. We further apply a pretrained GAN-based video upsampler [43] to super-resolve our generated videos 4 times, resulting a final 360° video of 1416x2832 resolution.
3.3 Reconstruction from “Generated” Captures
Given the generated 360° capture video with 128 panoramic frames, we reconstruct a 3D Gaussian scene. For each panoramic frame, we randomly perform 3 perspective crops with a field of view of 120° and resolution of 512$\times$512, resulting in 384 perspective images for pose estimation and reconstruction. Similar to standard scene reconstruction, we calibrate the camera poses of these perspective images using COLMAP [36]. The high average matching ratio $94.6\%$ from COLMAP estimation demonstrates the strong view consistency of the generated videos from our text-to-360° video generator. We adopt Long-LRM, the state-of-the-art feed-forward 3DGS reconstructor [56] to reconstruct the final 3D scene using the estimated camera poses and generated views. Using Long-LRM allows us to achieve fast reconstruction compared to an optimization-based approach. We randomly subsample 32 random views along the original camera path as the inputs since the Long-LRM has a maximum limit for input views. The calibrated camera poses from COLMAP are converted into Plücker rays as the inputs for long-LRM.
3.4 Implementation Details
We fine-tune a pretrained text-to-video diffusion model using mixed 360° image and video datasets. The fine-tuning is conducted on 32 NVIDIA A100 GPUs. 8 GPUs train on 360° images, while 24 generate 128-frame 360° videos. Of these, 12 GPUs produce videos at 24 fps, and 12 generate videos at 12 fps. The model was finetuned from a pretrained text-to-video checkpoint for 20,000 steps in around 200 hours. During inference, we generate videos at 12 fps to extend the traversal range while maintaining good temporal consistency comparable to 24 fps generation.
The inference of a 10.6 second video takes around 10 minutes with an unoptimized PyTorch runtime [30] using 50 diffusion steps, together with video upsampling process. The entire reconstruction takes around 5 minutes, including pose estimation and Long-LRM reconstruction. Our full 3D generation pipeline requires approximately 15 minutes to generate a detailed GS scene, with potential for further optimization to improve inference speed and efficiency for different components.
[…]
360DVD [42] 0.7643 3.5797 NA
DreamScene360 [55] 0.8116 4.0942 2.4125
Ours 0.7387 4.2444 2.8043
Table 1: We quantitatively evaluate our method against the state-of-the-art (SOTA) 360° video generation model [42] and the SOTA text-to-scene model [55], which generates 360° coverage scenes. CLIP distance and Q-Align metrics are computed on the generated 360° images or frames across 320 test text prompts. Additionally, Q-Align NVS assesses the quality of novel view synthesis.
A calm Zen rock garden in an ancient monastery courtyard.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A candlelit medieval cathedral crypt with carved statues.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A Celtic stone circle on rolling green hills at dawn.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A glacier-capped polar landscape with drifting icebergs.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A cramped submarine cafeteria with portholes showing ocean depths.
Refer to caption (a) Panorama Reference Refer to caption (b) Novel View 1 Refer to caption (c) Novel View 2 Refer to caption (d) Novel View 3 Refer to caption (e) Novel View 4
Figure 3: Scene generation results. We present our scene generation results by showing (1) the first 360° image from our generated panoramic video as reference, and (2) novel video renderings (Novel View 1-4) from different spatial locations and viewpoints, demonstrating that our approach effectively synthesizes a traversable, 360° viewable scene. More rendered videos and navigable Gaussian Splats scenes can be found in the supplementary documents.
A cottage in the woods where the walls are made of interwoven tree roots, and the floor is a carpet of moss.
DreamScene360
[55]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
A cottage surrounded by fields of wildflowers that change color with the seasons, glowing faintly at twilight.
DreamScene360
[55]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
(a) Panorama Reference
(b) Novel View 1
(c) Novel View 2
(d) Novel View 3
(e) Novel View 4
Figure 4: Comparison of scene generation results. We compare our method with DreamScene360 [55] for text-based scene generation, as both approaches use panoramas as an intermediate representation for generating 360° viewable scenes. The first column presents the panorama reference: for DreamScene360, it is a single generated panorama from a text-to-image model, whereas in our results, it is the first frame of our generated 360° panorama video. While the previous work can produce a 360° viewable scene, its range of motion is highly limited. When novel viewpoints (Novel View 1-4) deviate significantly from the initial position, their GS scene exhibits severe blurriness and artifacts. Besides, our method achieves better alignment with the input prompt and produces the correct global structure of the scene.
4 Results
We demonstrate that our pipeline achieves state-of-the-art quality in both 360° video generation (Sec. 4.1) and 3D scene generation (Sec. 4.2), offering significantly improved visual consistency and view coverage. More visual examples, including videos and Gaussian Splats with an interactive viewer, can be found in the supplementary material. To quantitatively evaluate our method, we prompt ChatGPT [28] to generate a list of 320 descriptions of 3D scenes. We then use CLIP distance [34] to assess prompt adherence and Q-Align [46] to evaluate visual quality.
Additionally, we perform ablation studies in Sec. 4.3 to evaluate our design choices, including mixed data training and masked loss, to ensure the generation of high-quality videos suitable for reconstruction.
4.1 Text-to-360°-Video Generation
Our generation model can generate temporally consistent 360° panoramic videos, exceeding the quality of the state-of-the-art panoramic video generation models. We compare our 360° video generation results with 360DVD [42] in Fig. 5. We showcase four evenly sampled frames comparing 360DVD and our 360° panorama video generation model. Our model generates videos that closely match the input text prompt and produce richer scene details compared to 360DVD. The higher average CLIP distance and Q-Align metric for scenes generated from the 320 prompts in Table 1 further demonstrate the superior prompt alignment and visual quality of our method.
Additionally, videos generated by our model cover a broader scene span compared to 360DVD, which exhibits only slight rotations and translations of the viewpoint as indicated in the marked green box (Fig. 5). More importantly, since the primary goal of 360° panoramic video synthesis is to generate a coherent 3D scene, the quality of 360DVD’s results prevent it from effectively accomplishing this task. Its outputs lack structural consistency across views, exhibiting continuous morphing and noticeable changes, which undermine the perception of a stable 3D environment.
To validate this statement, we run COLMAP on their generated panoramas using a similar cropping and pose estimation procedure as described in Sec. 3.3. We find that the failure rate of COLMAP reaches $89.7\%$, and even in cases where COLMAP succeeds, the match rate remains very low, with only $41.8\%$ images being successfully matched, making their results unsuitable for 3D reconstruction. In contrast, our model preserves view-consistent local structures across frames. We do not encounter any failure cases when running COLMAP, and the average image match rate is as high as $94.6\%$ across our test prompts. This is the key reason we achieve a well-reconstructed 3D scene.
Ours DreamScene360 [55] LucidDreamer [10]
View Coverage 360° 360° Wide-angle front view
Movable Range Large Very Limited Limited
Table 2: Our method achieves state-of-the-art view coverage and traversable range compared to previous 3D generation methods.
4.2 Text-to-Scene Generation
As shown in Fig. 3, by chaining our 360° video generator with the 3D reconstruction process, we enable the creation of a diverse set of 3D scenes, allowing users to navigate the environment from novel viewpoints and spatial locations. More results can be found in the supplementary material.
In Fig. 4, we compare our results to the previous state-of-the-art text-to-scene method that can also create 360° virtual environment. However, as an approach that reconstructs a scene from only a single panorama, DreamScene360 cannot create a truly navigable scene. When synthesizing novel views (Novel View 1-4 in Fig. 4) from locations different from the initial position, the rendered images exhibit significant artifacts and blurriness. In Table 1, we compute the Q-align metrics on the rendered novel views (Q-align NVS), where our results achieve higher visual quality and less blurriness compared to previous work, demonstrating that our generated scene enables a more realistic navigation experience. Similarly, we also report the CLIP distance and Q-Align metrics on the generated panorama images, demonstrating that our method generates a 3D environment that more accurately responds to user requests while achieving superior visual quality, even at the initial position.
In addition to quantitative measurements, we highlight that our method is the first to achieve traversable and 360° viewable text-to-scene generation. As shown in Table 2, compared to previous work, our method achieves a complete 360° viewing area while also allowing users to walk inside the generated 3D scene for a large range.
4.3 Ablation Study
Training a high-quality 360° video generator within our 3D generative pipeline is a key factor in the success of our method, relying on the use of mixed training data and the masked diffusion loss. We conduct ablation studies on these design choices to demonstrate their effectiveness.
4.3.1 Mixed Training Data
We conduct an experiment by excluding the additional image panorama data. As shown in Fig. 6, incorporating image data significantly enhances the generalization capability of the video generator across a wide variety of scenes, and improves visual quality. Notably, due to the masked diffusion loss training, ground-truth pixels in the bottom regions of the panoramic videos are unavailable. The image data compensate for this limitation, enabling the model to intelligently auto-inpaint the bottom regions, resulting in a natural and realistic appearance.
4.3.2 Masked Diffusion Loss
We also train a model using the full diffusion loss applied to all pixels. The generated videos in this case reveal the photographer and the camera pod, which consistently appear in the 360° video training data (Fig. 7). This obstructs a significant portion of the generated capture videos, making comprehensive scene reconstruction challenging. By applying the masked diffusion loss and incorporating the aforementioned image panorama data, we guide the model to generate clean 360° videos that can be directly used for perspective cropping and reconstruction.
A cozy cabin library overlooking a vast snowy canyon.
360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
Ours
A coastal castle perched on a cliff under a dramatic sunset.
360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Figure 5: Comparison of 360° video generation results. We present a comparison of 360° panoramic video generation results against 360DVD [42]. Our generated videos exhibit significantly better visual quality and prompt alignment. We manually mark an anchor region with a green box in the 360° image, highlighting the substantially longer traversal range of our generated videos.
A cozy cottage surrounded by wildflowers that glow in the moonlight, nestled deep in an enchanted forest.
w/o Image Data Refer to caption Refer to caption
w/ Image Data
A serene desert where the sand shifts in swirling patterns, reflecting the colors of the sunset.
w/o Image Data Refer to caption Refer to caption
w/ Image Data
Figure 6: We verify that incorporating additional panorama image data helps the model generalize to diverse environments. Without the image data, the generated 360° videos not only exhibit degraded visual appearance but also display messy and incorrect content at the bottom of the video. This issue arises because these pixels lack supervision due to the masked loss.
A garden where every flower blooms in perfect circles, their petals reflecting the colors of the stars above.
Baseline Refer to caption Refer to caption
w/ Masked Loss
A room where the ceiling is made of glass, revealing the swirling, ever-changing night sky beyond.
Baseline Refer to caption Refer to caption
w/ Masked Loss
Figure 7: We conduct an ablation study to evaluate the impact of the masked diffusion loss. This loss function effectively guides the model to generate only the pixels of the scene, eliminating the moving photographer present in the original video training data. As a result, it enables the generation of a complete 360° scene capture for reconstruction.
5 Conclusion
In conclusion, we presented WorldPrompter, the first generative pipeline capable of producing a navigable 3D scenes from text prompts. By leveraging 360° panoramic video generation as an intermediate representation and a fast 3D reconstruction model, WorldPrompter enables the synthesis of immersive environments with comprehensive view coverage and high visual quality. This approach bridges the gap between text-to-3D generation and full scene synthesis.
Although WorldPrompter achieves state-of-the-art performance, generating longer panoramic videos, spanning minutes rather than seconds, remains an open challenge. Consistent content creation over extended sequences is crucial for scaling up to larger, more detailed 3D scenes. Addressing this challenge would further expand applications in virtual worlds and beyond, paving the way for future advancements in scene-level generative models.